OpenStudy (anonymous):

Devin wants to have equal amounts of room on all four sides of his ping pong table. A 5ft by 9ft table is placed in the center of a rectangular room that has an area of 221 square feet. What is the width of the clear space Devin will have on all sides of his table.

hero (hero):

(2x+9)(2x+5)=176

hero (hero):

Area of Room = 221 Area of Table = 45 221 - 45 = 176

hero (hero):

Solve for x

hero (hero):

Do you know if the answer is supposed to be a whole number or not? I get about 3.2 feet on all sides

OpenStudy (anonymous):

Im doing it

hero (hero):

Okay

OpenStudy (anonymous):

i got 4

hero (hero):

I see, my formula was wrong You're supposed to either use (2x+5)(2x+9) = 221 or (2x-5)(2x-9) = 176

hero (hero):

Yeah, your way of doing it is correct x = 4